Solve the following linear equations. 
n/2 – 3n/4 + 5n/6 = 21

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

n/2–3n/4+5n/6=21
⇒(6n–9n + 10n)/12 = 21
⇒7n/12=21
⇒7n = 21 × 12
⇒n = 252/7
⇒n = 36

Thanks